c=100+0.9yd
i=100
g=200
t=160
nx= -200
y=?
1
Expert's answer
2014-05-28T08:31:43-0400
We can write an equation:

Y= C + I + G + NX, where

C- Consumer spending is a personal consumer expenditures for all citizens that includes the costs of durable goods and the
current consumption and the services. This is an expense for food, clothing,
housing, cultural products, household products etc.
I - Investment costs cover the basic types of private investment in production by firms and entrepreneurs. These costs are used for the ultimate purchase of machinery and equipment, industrial buildings, the increase in reserves of inputs, housing etc.
G - The government spending is government purchases of goods and services. This group includes the costs of federal, national and local government purchases of the companies’ final product for public use. The government spending does not include transfer payments.
NX - Net exports shows how foreign spending on goods and services produced in the country exceed the country costs for the purchase of goods and services produced in other countries.

So,we can write:
Y= 100+0.9yd + I + G + NX , where
yd- disposable income.
yd= Y – T
Y= 100+0.9*( Y – T) + I + G + NX
Y= 100+0.9*( Y – 160) + 100 + 200 – 200
Y= 56 + 0.9Y
0.1Y= 56
Y= 560